You must give a dog 250mg of a medication that comes in tablets labeled 100mg each. How many tablets will you give? 2.5 A 20kg dog requires 50mg/kg of a medication that comes in 2g tablets. How many tablets will you give? ________ A cat requires 50mg of a liquid medication that has a strength of 150mg/ml. How much will you give? ______
Answer #1

Ans) Dose: 250 mg

Available: 100 mg tablet

100 mg = 1 tab

250 mg = x

x= 250/100 = 2.5 mg

20 kg dog

Dose: 50 mg/kg

50 × 20 = 1000 mg = 1 gm

So available is 2 gm so ½ tablet to jh

Available: 150 mg/ml

Dose= 50 mg

150 mg = 1 ml

50 mg = x

x= 50/150 = 0.33 ~ 0.3 ml